Skysong Co. invested $1,010,000 in Concord Co. for 25% of its outstanding stock. Concord Co. pays out 30% of net income in dividends each year. Use the information in the following T-account for the investment in Concord to answer the following questions. Investment in Concord Co. 1,010,000 99,000 29,700 (a) How much was Skysong Co.’s share of Concord Co.’s net income for the year? Net income $ (b) What was Concord Co.’s total net income for the year? Total net income $ (c) What was Concord Co.’s total dividends for the year? Total Dividends $ (d) How much was Skysong Co.’s share of Concord Co.’s dividends for the year? Dividends $
Answer #1

a) Share in net income = $99000

b) Total net income = 99000*100/25 = $396000

c) Total dividend = 29700*100/25 = 118800

d) Share in dividend = $29700